Consider a two countries, Portugal and England, that produce two goods, wine and cheese, with only one factor of production, Labor. In England, one unit of labor can produce 2 units of wine or 1 unit of cheese. In Portugal, one unit of labor can produce 3 units of wine or 1/2 of cheese. There are 100 units of labor in Portugal, and 100 in England. Countries share the same tastes, and there is perfect competition.

1) Fill in the following table:

England

Portugal

Wine

Cheese

Labor

PW/PC

PC /PW

  1. Which country has a an absolute advantage in wine? In cheese? Which country has a comparative advantage in wine? In cheese?
  2. Draw the Production Possibilities Frontier for each country.

Remember that the production possibilities frontier is the combination of wine and cheese that can be producedin each country, given resources (in our case, given the amount of labor).

  1. Describethe autarkyequilibrium in each country. (Represent graphically if you wish.)

Remember that for equilibrium we want to know how much each country is producing and consuming of each good (Hint: you will NOT be able to give a precise numerical answer.) AND the relative price of goods.

a)Suppose now that preferences are such that consumers in both countries (remember, tastes are the same) always want to consume twice as much wine as cheese. What would the autarky equilibrium be? (Now you can give me a precise number!)

  1. Suppose the two countries are allowed to trade. What will be the range of relative prices of cheese for which there will be trade? What will be the range of relative price of wine for which there will be trade?

(Notethat when you read “relative price of cheese” this is means “price of cheese relative to wine” or PC/PW)

  1. Suppose the world relative price of cheese is 3 . (Ignore the preferences I gave in question 4a, which were for illustrative purposes only).
    1. Will the countries trade? If so, describe the pattern of trade.
    2. At this price, exactlyhow much wine and cheese will Portugal produce? In particular, what can you tell me about specialization?
    3. Explain the gains from trade.
  2. Suppose the world relative price of cheese is 4. What changes with respect to quest. 6?

Does the pattern of trade change? Does the level of production change? Do the gains from trade change?

For the gains from trade, look carefully at how the Consumption Possibilities changes with the new price.

  1. Suppose the amount of labor in Portugal grew to 120. How would the PPF in Portugal change? Would Portugal’s comparative advantage change? Would the range of prices for which trade occurs change?
  2. Suppose now the world relative price of cheese is 2. Will there still be trade? Why or why not? (HINT: Once again, look at the gains from trade)
Answer #1

1) If England uses all labor to produce wine, then wine production will be 100*2 = 200 units and if all labors are used to produce cheese, then total cheese will be 100*1 = 100 units.

Similarly, If portugal uses all labor to produce wine, then wine production will be 100*3 = 300 units and if all labors are used to produce cheese, then total cheese will be 100*1/2 = 50 units.

Thus, the table will look as follows -

England Portugal
Wine 200 300
Cheese 100 50
Labor 100 100
PW/PC 100/200=0.5 50/300 = 0.167
PC/PW 200/100=2 300/50=6

2) Absolute advantage occurs in country which produces more units with given resources. As you can see, Portugal can produce more of wine and England can produce more of cheese with 100 labors as compared to each other, then Portugal has absolute advantage in producing wine and England has absolute advantage in producing cheese.

Comparative advantage occurs when a country has lesser opportunity cost in production of a good than the other country. It is same as price of good in terms of other good (such as PW/PC and PC/PW). As you can see PW/PC (price of wiine in terms of cloth) is lesser of Portugal, then Portugal has comparative advantage in producing wine. And PC/PW is less of England, thus England has comparative advantage in producing cheese.

3) Use the given coordinates of cheese and wine for both country to draw PPF curves-

England (200,0) and (0,100)

Portugal (300,0) and (0,50)

Wine is placed at X axis and cheese is placed at Y axis

Cheese 120 100 PPF england 80 60 40 20 PPF Portugal 50 100 150 200 250 300 350 wine

4) In autarky equilibrium, both countries will produce any combination given at their respective PPF curves. The relative price equals PW/PC and PC/PW for each country given in the above table.